Determine the statical determinacy of the truss below and find the forces in all members of the truss due to the loading shown below. State if the member is in tension or compression.

3m EG. 15KN 15KN 4 panels at 4 m = 16 m
Answer #1

The detailed solution is given below. Number of members(m) = 13 Number of joints (g) = 8 hos 21-3 = 2x8-3 3 13 am So, (m=2; 3) The truss is stable and statically d using Method of Joints Initially we assume all the members are, an tension. If the value obtained as a negative valued, thenToint H FAF 0-36.890 RH FAG T 11:25kn *£Fy=0 => Fue Sino+11.25 = 0 => FHF = -18.75kNT A E Fx = 0 → FHF Coso + FAG=0 → FHG = 1Efyzo FFHCosa + FFE Cosa so * (FFE = - FHF = 18.75 kN] * EFxzo > FAF Sind- FFDFFE Sind=0 > FED = -18.75 sud - 18.75sma → FED

Please give THUMBS UP if you have liked it.